OpenStudy (anonymous):

Solve (x+9)/(x+8)=(x-7)/(x-6)

OpenStudy (anonymous):

find a common denominator first

OpenStudy (luigi0210):

Cross multiply: \[\frac{x+9}{x+8}=\frac{x-7}{x-6}\] So it's: \[(x+9)(x-6)=(x-7)(x+8)\] Distribute and solve now. It's a bit more complicated this way but it works as well.

OpenStudy (anonymous):

Thanks!

OpenStudy (luigi0210):

Yup

OpenStudy (luigi0210):

No medal? ;-;